THE ULTRAVIOLET SPECTROSCOPY OF PHENYLCYCLOPENTENE AND PHENYLCYCLOPENTADI-ENE

\underline{JOSH J. NEWBY}, CHING-PING LIU, CHRISTIAN MÜLLER and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

Recent studies of the effluents of an electric discharge of benzene, 1,3-butadiene, and vinylacetylene independently observed a resonant two-photon ionization (R2PI) spectrum with mass \frac{m}{z} = 142 whose carrier was unidentified. Based on the similarity of its low frequency vibronic structure to that of α-methylstyrene, we have synthesized phenylcyclopentadiene, recorded its R2PI spectrum, and found it to be the carrier of the discharge-produced \frac{m}{z} = 142 spectrum. The molecule exists in two isomeric forms: phenylcyclopentadi-1,3-ene and phenylcyclopentadi-1,4-ene. UV-UV holeburning spectroscopy was used to record the UV spectra of the two possible isomers. The suspected origin transitions of the two isomers are found at 31738 and 31687 cm^{−1}, respectively. Using similar reasoning, an analogous spectrum observed in a butadiene discharge at m/z 144 is ascribed to phenylcyclopentene, with an S_{0}-S_{1} origin at 34646 cm^{−1}. In addition to recording the R2PI spectrum, LIF excitation and single vibronic level dispersed fluorescence scans of phenylcyclopentene and phenylcyclopentadiene have been recorded. In phenylcyclopentene, a progression involving the out-of-plane torsion of the two rings was identified, with a vibrational frequency of 44 cm^{−1} in S_{0} and 72 cm^{−1} in S_{1}. An analysis of this torsional vibronic structure and its underlying torsional potential will be described and compared with ab initio predictions.